OpenStudy (anonymous):

Solve: 64x^6-65x^3=-1

OpenStudy (anonymous):

solve \[64z^2-65z+1=0\] the replace z by \(x^3\)

OpenStudy (anonymous):

get \[(64z-1)(z+1)=0\] \[z=-1,z=\frac{1}{64}\] then solve \[x^3=-1,x^3=\frac{1}{64}\]

OpenStudy (anonymous):

ok that should give me 6 answers correct

OpenStudy (anonymous):

i think satellite has made a little mistake. \[64z^2-65z+1=0\] after that it should be \[(64z-1)(z-1)=0\] then \[x^3=1, \frac{1}{64}\] I think there are only 2 answers which are 1 and 1/4. P.s. if there are mistakes, please tell me.

OpenStudy (anonymous):

but the original problem is x^6